Given the same problem in #2. Assume that the capacitors are connected in parallel.

a) Find the equivalent capacitance of the circuit.

b) Solve for the charge across each capacitor.



1
Expert's answer
2022-03-21T12:42:08-0400

Given: C1 = 2.0 μF, C2 = 4.0 μF, C3 = 5.0 μF, C4 = 7.0 μF, V= 9.0 V.

The capacitors are connected in parallel.


a) Find the equivalent capacitance of the circuit.

"C=C_1+C_2+C_3+C_4\\
=2.0+4.0+5.0+7.0=18\:\rm\mu F"


b) Solve for the potential difference across each capacitor.

"V_1=V_2=V_3=V_4=V=9\:\rm V"